### v1

# The HITRAN2024 methane update

Abstract
Spectroscopic parameters of methane from many different studies were gathered to improve the HITRAN database towards its 2024 version. After a validation process using high-resolution FTS and CRDS spectra, about 60,000 lines of the four most abundant isotopologues were replaced from the dyad to the triacontad regions. These changes amount to $56,000$ transition wavenumbers, $16,000$ line intensities, $30,000$ pressure-broadening half-widths, and 3500 assignments. 40,000 new lines were added, extending the database from 12,000 cm-1 up to 14,000 cm-1, and covering some gaps. A greater focus was brought on the pentad, octad, and tetradecad regions, targeted by several remote instruments. In these regions, comparisons of spectral fits from multiple line lists were performed, taking only the parameters that provide best fit for each line. In the $\nu_{3}$ band, in addition to replacing the previous values, gathered speed-independent pressure broadenings of 12CH4 were used to fit Padé-approximant functions. These functions then replaced any outdated experimental data in $\nu_{3}$, missing data in the new lines, as well as the values that were determined to be outside their physical boundaries. The CH3D broadenings were replaced in the same manner, for missing and low or high values, using a semi-empirical formula instead.
keywords:
HITRAN; Spectroscopic database; Molecular spectroscopy; Spectroscopic line parameters; Spectral line shape; Line broadening; Line shift
journal: J. Quant. Spectrosc. Radiat. Transfer
1 Introduction
The spectroscopy of methane has a fundamental role in the observations and radiative transfer modeling of gaseous media, especially atmospheres of Earth [chanmiller_methane_2024], solar system planets and moons [10.1006/icar.1994.1139, 10.1016/j.icarus.2005.02.004], exoplanets (e.g. HD 102195b [10.1051/0004-6361/201834615]), brown dwarfs (e.g. HD 102195b [10.1093/MNRAS/STAC1412]), and combustion environments [10.1364/AO.35.004026]. The recent development of remote observation tools such as those on board Sentinel-5 [lorente_methane_2021], GOSAT, GOSAT-2 [imasu_greenhouse_2023], and MethaneSAT [chanmiller_methane_2024] has pushed the demand for higher quality parameters at many temperatures and pressures, in the presence of various other species, provided by the HITRAN database [gordon_hitran2020_2022]. Methane is frequently measured in planetary and stellar atmospheres, where its study plays an important role in understanding the formation and composition of these astronomical objects. Important methane-specific data such as its presence, concentration, or partial pressure would not be retrievable without models that incorporate accurate spectroscopic parameters.
The octad (3500–5000 cm-1) and tetradecad regions (5000–6250 cm-1) in particular are crucial for satellite missions [lorente_methane_2021, malina_consistency_2022, chesnokova_estimation_2020], not only for its principal isotopologue but also for 13CH4 and CH3D, which are used to determine 13C/12C and D/H ratios. The octad has been improved in previous editions of HITRAN in the window observed by TROPOMI (4190–4340 cm-1). However, at higher wavenumbers, spectral parameters were of lower accuracy, incomplete, or both. Current and planned CH4 satellite sensors strive for  0.5% precision in the column (Ref. [jacob_quantifying_2022]; Table 1) for area flux mapping purposes, which requires similar accuracies for the intensities and widths and is not achieved in many regions. Even in the TROPOMI window, the HITRAN uncertainty code only reaches 4 (i.e., $0.0001\leq\Delta\nu<0.001$ cm-1 for line positions and $10\%\leq\frac{\Delta S_{ij}}{S_{ij}}<20\%$ for intensities). Above, and including the tetradecad region, other problems arise from the lack of assignment and/or value for the lower state energy ($E^{\prime\prime}$). Given the temperature range, these issues are limited in the Earth’s atmosphere, but make accurate spectral modeling impossible at more extreme temperatures (for example, low temperatures on Titan or high temperatures on hot-Jupiters and brown dwarfs). Furthermore, the origin of some of the parameters present in HITRAN2020 is not well documented, especially for the pressure-induced data. This can cause additional problems, preventing comparisons or adjustments in the case of models.
The focus of the HITRAN2024 update was on transition wavenumbers, intensities, line-shape parameters due to pressure of air, and assignments. Significant progress has been made between 900 and 14000 cm-1 [kefala_empirical_2024, nikitin_improved_2024, rodina_improved_2021, nikitin_preliminary_2013, starikova_assignment_2024, malarich_dual_2021, nikitin_first_2024, campargue_high_2023] for the three most abundant isotopologues, 12CH4, 13CH4, and 12CH3D, allowing for an update to the transition wavenumber, intensity, and assignments from the HITRAN2020 database using various recorded spectra for validation. These line lists also cover missing portions of the infrared spectrum that were previously challenging to obtain, as well as some in-between lines that were missing. Similar efforts have been put into obtaining good quality pressure-induced line-shape parameters [antony_n2_2008, bertin_co2_2024, clark_difference_2004, devi_self-_2015, devi_spectral_2016, devi_multispectrum_2018, es-sebbar_intensities_2014, es-sebbar_linestrengths_2021, farji_airinduced_2021, gabard_line_2010, gharib-nezhad_h2induced_2019, ghysels_spectroscopy_2011, ghysels_temperature_2014, grigoriev_estimation_2001, lyulin_measurements_2014, ma_temperature_2016, manne_determination_2017, mondelain_line_2005, mondelain_measurement_2007, pine_speeddependent_2019, pine_speeddependent_2000, pine_multispectrum_2003, richard_self_2023, dudaryonok_semiempirical_2018], with the differences being that we considered older work, with the intention of replacing the parameters for more lines than those in Section 2 to improve traceability. To this end, a combination of good quality measurements, with an empirical Padé-approximants model based on these measurements and semi-empirical models to fill the gaps was used. These results provide the the necessary foundation for the inclusion of first-order and full line-mixing with speed-dependent profiles.
2 Additions and updates
The newest versions of MARVELized-ExoMol [yurchenko_exomol_2024], TheoReTS [10.3847/1538-4357/aa8909], HITEMP [10.3847/1538-4365/ab7a1a] and MeCaSDa [MeCaSDa_2013], which is partly accessible via VAMDC portal [rixon_VAMDC_2011, albert_VAMDC_2020], line lists as well as lists from Nikitin et al. [nikitin_improved_2024, nikitin_preliminary_2013, nikitin_first_2024], Rodina et al. [rodina_improved_2021], Sung et al. [sung_new_2024], Starikova et al. [starikova_assignment_2024], Malarich et al. [malarich_dual_2021], and Campargue et al. [campargue_high_2023] were available for the HITRAN2024 update and are summarized in Sec. 2.1. These works were then validated against several high-quality, high-resolution FTS [2015ApJ...813...12H, bertin_co2_2024, birk_esa_2017, birk_measurement_2025] and CRDS spectra [birk_esa_2017] obtained under well-controlled laboratory conditions. The best parameters depending on the spectral region were selected and added to the HITRAN2024 database.
Much of the focus was brought to the pentad, octad, and tetradecad regions, but many other spectral regions have also been updated for HITRAN2024. Changes from these lists were applied almost exclusively for the transition wavenumbers, intensities, and energy levels, pressure-induced parameters of most of the available line lists were missing or approximate. Updates on these parameters are detailed in Sec. 3. The newly added lines, which were formerly missing in HITRAN, were added as is, but their pressure broadenings and shifts were replaced as well.
2.1 New line lists at disposal
This section summarizes recent studies for methane infrared spectroscopy. Each line list was subjected to a validation process to determine the parameters to include in the HITRAN2024 update. The specific additions from these works are described in Sec. 2.2.
2.1.1 ExoMol
The ExoMol project [tennyson_2024_2024] has produced a number of methane line lists of increasing completeness and accuracy designed for use at high temperatures [yurchenko_exomol_2014, yurchenko_hybrid_2017, yurchenko_exomol_2024]. The intensities in this list are calculated variationally. The line positions in the most recent of these [yurchenko_exomol_2024], known as the MM line list, was based on a comprehensive MARVEL (measured active vibration-rotation energy levels) study of 12CH4 by Kefala et al [kefala_empirical_2024]. This MARVEL study extracted 82173 rovibrational and rotational transitions from 96 literature sources containing high-resolution studies of 12CH4 spectra; the study identified a further 42 sources containing unassigned or partially assigned methane spectra which could not be included in the MARVEL study. An analysis of these sources is currently in progress and has so far led to the assignment of almost 10000 further lines, which extend the coverage of the previous study in terms of coverage of both rotational and vibrational states. This work can enable updates to future releases of both HITRAN and HITEMP. A HITRAN-format 296 K version of the MM line list generated using pyExoCross [Zhang_pyExoCross_2024] was used in this study.
2.1.2 TheoReTS
The TheoReTS database [rey_theorets_2016] provides line lists and quasi-continuum theoretical calculations for low and high temperatures from ab initio potential energy surfaces (PES) and dipole moment surfaces (DMS) together with spectra simulation tools. It was recently updated in [10.3847/1538-4357/aa8909] [rey_new_2018]. Variational line position calculations by Rey et al. have been refined using available empirical energy levels deduced from effective spectroscopic models. The TheoReTS predictions have been experimentally validated via high-temperature laser experiments by Ghysels et al. [ghysels_temp_2018] and by Wong et al. [Wong_atlas_2019] for their methane cross sections, as well as in dual frequency comb experiments and double resonance experiments by Malarich et al. [malarich_dual_2021] and Foltynowicz et al. [foltynowicz_2021] respectively. They were also used by Rey et al. [rey_new_2018] to model methane absorption at Titan, as observed by VIMS and by DISR instruments onboard the Cassini and Huygens space missions, and for the interpretation of non-LTE experiments in a hypersonic jet flow by Dudas et al. [dudas_2023]. In the context of the present HITRAN2024 update, TheoReTS predictions were used to create the HITEMP methane line list [10.3847/1538-4365/ab7a1a], while the extended assignments of 12CH4, 13CH4 and 12CH3D experimental spectra were carried out with improved line lists as described in the following sections. In recent multi-laboratory line intensity measurements, Reed et al. [reed_multilaboratory_2025] reported permille-level uncertainties of the R(0) to R(10) manifolds in the CH4 $2\nu_{3}$ band. These results had an average agreement of about 0.05 % with TheoReTS calculations which used the ab initio DMS of ref. [NIKITIN_dipole_intens_2017].
2.1.3 HITEMP
The HITEMP spectroscopic database [10.1016/j.jqsrt.2010.05.001] contains line lists for eight molecules, with the methane line list fully described in Hargreaves et al. [10.3847/1538-4365/ab7a1a]. The HITEMP list was constructed by combining the empirically corrected ab initio line list for 12CH4 from TheoReTS [10.3847/1538-4357/aa8909] with the HITRAN2016 line list [gordon_hitran2016_2017], which was the contemporary HITRAN list at the time. HITEMP CH4 spans 0-13,400 cm-1 and is suitable for simulating the high-temperature spectrum of methane up to 2000 K. In addition, the HITEMP list includes “effective” lines that are able to reproduce the continuum-like features provided by temperature-dependent lists of weak lines (i.e., so-called super-lines) in TheoReTS. Readers are referred to Hargreaves et al. [10.3847/1538-4365/ab7a1a] for a full description of the “effective” line methodology for methane.
Merging HITRAN2016 with the TheoReTS line list required consistency with assignments, as well as positions and intensities to ensure uniqueness of transitions. Below 6230  cm-1, the HITRAN2016 line list contained both upper-state and lower-state assignments for vibration and rotation quanta; however, above this point, the assignments were limited. The assignments of the TheoReTS list were also incomplete, which enabled lines to be matched only by their positions, intensities, lower state energies, and rotation number ($J^{\prime\prime}$). Consequently, approximately 81,000 ($\sim$50%) of 12CH4 lines in HITRAN2016 replaced corresponding lines of TheoReTS. Furthermore, there were no equivalent high temperature line lists for isotopologues of methane (13CH4, 12CH3D, 13CH3D) and these were imported into HITEMP directly from HITRAN2016. The resultant HITEMP line list contains $\sim$32 million lines, of which approximately 5 million are effective lines and for which the largest contribution of lines originates from the 12CH4 line list of TheoReTS.
2.1.4 MeCaSDa
CaSDa (Calculated Spectroscopic Databases) [RICHARD2024109127] is a project that presently gathers ten databases of line lists for ten highly symmetrical molecules including methane. These lists are built upon the analysis of high-resolution spectra (microwave, infrared, and Raman) using effective Hamiltonian and dipole moment operators written with a tensorial formalism [BOUDON2004620]. The databases also contain the full description of eigenvectors. Within this project, the MeCaSDa database was built for 12CH4 and 13CH4. The last update of MeCaSDa was performed in 2020 [RICHARD2020107096], following the results of a global fit described in Ref. [10.1063/1.5023331].
2.1.5 IAO FTIR measurements 900–1050 cm-1
Four spectra of methane samples at natural isotopic abundance were recorded between 900–1050 cm-1 at the Institute of Atmospheric Optics (IAO) in Tomsk [nikitin_improved_2024]. Nonlinear least-squares curve fitting procedures were used to retrieve line positions and intensities, which were subsequently analyzed using effective Hamiltonian and effective dipole moment models. The enhanced absorption at long paths allowed the measurement of new transitions: up to $J=28$ for cold bands and up to $J=20$ for hot bands. The new experimental list contains line positions and intensities for 2570 lines, with quantum assignments made for 1244 of these lines for the main isotopologue 12CH4. Comparisons of theoretical absorption simulations with experimental spectra revealed a considerable improvement compared to the HITRAN2020 database. All of the 1244 assigned line positions were fit with an average RMS standard deviation equal to 0.00065 cm-1.
2.1.6 Kitt Peak & JPL FTS measurements of 13CH4 4000–4700 cm-1
Multiple sets of 13CH4 line lists, either empirical or model predictions, were available from various sources, e.g., HITRAN 2020 [gordon_hitran2020_2022, brown_measurements_2016], JPL GFIT package (“atm.161”) [toon_spectrometric_2021], and Kitt Peak FTS (291 K) and JPl-FTS (80 K). More importantly, comprehensive updates to the previous work by Brown et al. [brown_measurements_2016] became available through the effective Hamiltonian modeling extended to tetradecad region for hot band transitions. All of the line lists have been evaluated and cross-compared against high-resolution laboratory spectra obtained at 291 K and 80 K, by which their quantum identifications and lower state energy values have been determined as presented in Sung et al. [sung_new_2024]
[…]
Table 1: Statistics on the line parameters in comparison to the existing spectroscopy databases and model calculations for 13CH4.
13CH4 line list Range Min. ($S_{ij}$) lines listed/ assigned
HITRAN2020 4000–4700 cm-1 $1\times 10^{-26}$ cm-1atm-1 761 / 693
Coy, Bourssard, Toon, Sung [sung_new_2024] 4000–4700 cm-1 $1\times 10^{-26}$ cm-1atm-1 7103 / 6585
A preliminary version of the composite line list has been implemented, followed by a final version to be available on-line, in which more quantum assignments will be available. Table 1 presents statistics on the improvements (both additions and updates inclusive) as reported by Sung et al. [sung_new_2024], showing an almost ten-fold increase in the total number of line list entries in the 4000–4700 cm-1 wavenumber interval for 13CH4, when compared to the HITRAN2020.
Figure 1: Comparison of synthetic spectra generated with HITRAN2020 (in pink) and the updated composite line list, our update (in red) with high-resolution Kitt Peak FTS spectrum at 291 K (Left, in black) and JPL-FTS spectrum at 80 K (Right, in blue), showing better performance and spectral coverage by the new composite line list. Note that the synthetic spectra were computed under the same conditions as the corresponding laboratory spectra. Any H2O (not presented) and 12CH4 features (in green) were identified during the line-by-line matching, as presented for 12CH4 features.
Finally, Fig. 1 presents a graphical summary of the updates, showing synthetic spectra generated with the HITRAN2020 line list and the composite line list reported in this work compared to two high-resolution laboratory FTIR spectra, respectively at 291 K( Left panel) and 80 K (Right panel). The synthetic spectra at 291 K and 80 K, respectively, were computed under the same conditions as the corresponding laboratory spectra. It is shown that the composite line list covers much wider spectral features across the Octad region as well as provides a better representation of the observed spectra at both temperatures thanks to the lower-state energy updates.
2.1.7 IAO analyses of FTS measurements 4970–6200 cm-1
Figure 2: An overview of the 13CH4 tetradecad line lists available in HITRAN2024. Newly added line parameters for 13CH4 are highlighted in magenta. The TheoReTS ab initio based variational list [rey_new_2018] with intensity cut-off $1\times 10^{-29}$ cm/molecule (grey dots) is plotted as a background overlapping with the colored data
Starikova et al. [starikova_assignment_2024] reported the first assignments of 13CH4 absorption bands belonging to the lower part of the tetradecad. This work focused on the analysis of a high-resolution Fourier-transform infrared spectrum of 13C-enriched methane in the 4970–5300 cm-1 range (Fig. 2). The spectrum was recorded over a broad range (4970–6200 cm-1) at 298 K using a 20.94 m path cryogenic Herriott cell vacuum-coupled to a Fourier transform spectrometer, Bruker IFS-125HR, at the Jet Propulsion Laboratory. The retrievals of line intensities from the experimental spectrum were carried out using the SpectraPlot [nikitin_visualization_2011] program, with the Voigt line profile at fixed values $\gamma_{\mathrm{self}}=0.08\ \mathrm{cm}^{-1}\mathrm{atm}^{-1}$, $n_{\mathrm{self}}=0.85$, and $\delta_{\mathrm{self}}=-0.018\ \mathrm{cm}^{-1}\mathrm{atm}^{-1}$. The uncertainties in the intensities of isolated lines were estimated to be about 10–15%, while those of blended lines were about 20–40%. Initial spectral assignments were obtained using the effective Hamiltonian model and effective dipole transition moment parameters, both derived from the previous work on cold 13CH4 spectra recorded at 80 K [starikova_13ch4_2018, starikova_assignment_2019, starikova_assignment_2016]. The published experimental list contains 1642 lines, of which 1600 transitions were assigned to four vibrational bands $4\nu_{4}$, $\nu_{2}+3\nu_{4}$, $\nu_{3}+2\nu_{4}$ and $\nu_{1}+2\nu_{4}$ of the tetradecad system, including 22 sub-bands with rotational quantum numbers up to $J_{\mathrm{max}}=16$. The parameters of the effective Hamiltonian and the dipole transition moment allowed the description of 1548 fitted experimental line positions with an rms deviation of $1.7\times 10^{-3}\mathrm{cm}^{-1}$, and the reproduction of a set of 674 measured line intensities with an rms deviation of 6.8%.
Figure 3: Comparison between the experimental spectrum of 13CH4 at 298 K (lower panel) with simulation based on the effective Hamiltonian and dipole transition moment models (middle panel) and ab initio TheoReTS [rey_new_2018] calculations (upper panel) in the 5000–5220 cm-1 tetradecad region. The blown-up part at the bottom corresponds to the narrow spectral interval around 5178 cm-1.
A comparison of the observed spectrum with the simulations using calculated effective parameters and the TheoReTS [rey_new_2018] line list, depicted in Fig. 3, shows the excellent quality of the calculations for this molecule as well as the improvements due to the reduction of experimental data.
2.1.8 NIST, DLR & USTC measurements 5970–6250 cm-1
The tetradecad spectral region has been the subject of extensive study, including the WKLMC empirical lineline [campargue_2013], the GOSAT 2014 line list [nikitin_gosat-2014_2015] (which was the basis of both HITRAN2016 and HITRAN2020 for this spectral region), the widely used empirical line list of Devi 2015 [devi_self-_2015], and the Nikitin 2017 analysis of WKLMC [nikitin_analysis_2017]. Significant differences exist between these line lists, complicated by correlation between parameters in these dense and complex spectra and the difficulty in identifying and assigning weaker transitions. These complications lead to line lists that trade parameter accuracy for self-consistency, thus impeding efforts to refine line parameters.
Recently, this problem has been examined in the methane tetradecad via three separate sets of experiments focused on the strong $2\nu_{3}$ band of 12CH4, which represents $\approx{67\%}$ of the total summed intensity of the R-branch spectral region from 6010 cm-1 to 6120 cm-1. Each experiment was tailored for optimal determination of particular spectroscopic parameters: saturation spectroscopy for transition frequencies, linear absorption spectroscopy in the Doppler-broadened domain for absolute transition intensities and ratios, and line shape and line mixing parameters determined in the pressure broadened domain. Transition frequencies for 53 transitions in the $2\nu_{3}$ R-branch were determined with kHz-level uncertainty through saturation spectroscopy by Votava et al. [votava_2022]. These highly accurate values were subsequently used as constraints in the determination of molecular line intensities in the Doppler-broadened domain through independent experiments employing CRDS at NIST and FTS at DLR as reported by Reed et al. [reed_multilaboratory_2025], which additionally replaced $\approx 2600$ weaker non-$2\nu_{3}$ lines from the Nikitin 2017 line list. Permille-level uncertainties were demonstrated (evaluated by relative agreement between the two independent measurements) for 64 strong transitions, covering the $2\nu_{3}$ R-branch from $m=$ (1–11). These line intensities were employed by Yin et al.[yin_2025] in further CRDS studies involving elevated pressure of CH4 broadened by N2 to determine line shape and line mixing parameters.
[…]
Figure 4: Comparison of the intensities for the HITRAN2020 (left pane) and HITRAN2024 line lists (right pane) in the 12CH4 tetradecad, including the $2\nu_{3}$ band. The overlap set includes transition intensities determined at NIST (CRDS) and DLR (FTS) in the Doppler-broadened domain.
Significant improvements in fit quality over previously determined spectroscopic parameters are demonstrated from Doppler-broadened conditions to atmospherically relevant pressure-broadened cases. The R(10) manifold is especially highly blended and the accuracy of the line intensities is significantly limited by fitting correlation. Given the large observed differences between DLR and NIST in the R(10) manifold, the NIST parameters were selected and used to constrain the USTC lineshape determinations at elevated pressures. Consequently, the resulting parameters reported for R(10) should be considered an effective parameter set.
Agreement with TheoReTS [rey_theorets_2016] is excellent (at the several permille level) for the assigned $2\nu_{3}$ transition intensities, however, the lack of assignments in the experimental line list for many weaker (non-$2\nu_{3}$) transitions makes direct comparison difficult. Additionally, many weaker lines present in TheoReTS have been merged in the experimental line list as they cannot be separated due to finite signal-to-noise or fitting correlation, further complicating direct comparison.
2.1.9 LiPhy DAS measurements of 12CH3D 6204–6394 cm-1
In spite of the small magnitude of the CH3D/CH4 amount ratio, (nominally $5\times 10^{-4}$ at natural abundance), CH3D has a large relative contribution in the region (up to 70 % near 6300 cm-1 at 80 K) [wang_high_2010]) and contributes greatly to the absorption in the 1.58 µm methane transparency window [lu_ch3d_2011]. Nevertheless, the $3\nu_{2}$ band is not isolated but superimposed on a congested spectrum of lines involving other weaker bands which have to be characterized for a proper description of the absorption in the region. As described in [brown_methane_2013], since its 2012 edition, the HITRAN list was extended to the 6204-6510 cm-1 region by including for the 6204–6394 cm-1 interval, the empirical line list constructed in Lu et al. [lu_ch3d_2011] using diode laser absorption spectroscopy (DAS) of highly enriched CH3D at 296 K. This list of about 5500 lines included about 175 rovibrational assignments and more than 2700 empirical values of the lower state energy, $E^{\prime\prime}_{\mathrm{emp}}$, obtained from the ratios of intensities measured at 296 K and 80 K (2T-method).
The update of CH3D in this spectral region is based on the recent work of Ben Fathallah et al. [benfathallah_ch3d_2024] which extends that of Lu et al. The region of the DAS recordings is increased to include the 6099–6530 cm-1 region (11189 lines) and about 4800 $E^{\prime\prime}_{\mathrm{emp}}$ values were obtained by the 2T-method. Relying on the position and intensity agreements with the TheoReTS variational line list [rey_accurate_2014, rey_theorets_2016] about 2890 transitions could be fully rovibrationally assigned to twenty bands, fifteen of them being newly reported. All the reported assignments were confirmed by Ground State Combination Difference (GSCD) relations i.e. all the upper state energies (about 1370 in total) have coinciding determinations through several transitions (up to 8). The intensity sum of the transitions assigned between 6190 and 6530 cm-1 represents 76.9 % of the total experimental intensities at 296 K.
2.1.10 PLDL Dual frequency comb 6770–7570 cm-1
Malarich et al. [malarich_dual_2021] present new broadband, high-resolution dual frequency comb spectroscopy (DCS) of pure methane up to 1000 K and 300 Torr in the methane icosad (6770–7570 cm-1). The dual comb spectrometer spans 870 cm-1 with 0.00667 cm-1 point spacing and a wavenumber uncertainty of 10-5 cm-1. Light from the spectrometer was passed twice through a 46-cm-long optical cell (92 cm total pathlength) placed in a uniform three-zone furnace.
Thirteen total spectra were analyzed at five temperatures from 296-1000 K and 18-300 Torr. Simulations in the icosad region with HITRAN2020 show good agreement with the measured spectra at room temperature, but exhibit significant mismatch at higher temperature. This is not surprising given that the empirically based HITRAN2020 icosad line list originated from spectra measured below 280 K (WKLMC [campargue_2013]). The new high-temperature spectra were used to update many HITRAN2020 lower-state energies, line positions and self shifts, self widths, and power-law temperature dependence exponents for self width and shift.
Much of the high-temperature discrepancy between HITRAN2020 and the DCS data was due to methane absorption features that had been given a default value of lower-state energy (such as $E"\textsubscript{HITRAN}=999$ cm-1). Using multispectrum fitting to the high-temperature spectra, 4283 lower-state energies are assigned which had previously been given this default value. Existing lower-state energies were also updated for 92 lines, and the database adds 293 high-temperature lines that improve the database performance above 300 K. The method for these updates is described in detail in [malarich_dual_2021]. Fig.8 in Ref. [malarich_dual_2021] shows the distribution of new lower-state energy values.
[…]
A similar approach was followed to determine band-wide average adjustments to self broadening (+7%), and to offer new power-law temperature scaling exponents of $n_{\gamma}=0.85$ and $n_{\delta}=0.58$. Altogether, these updates greatly reduce the model-measurement residuals for the icosad region across a large array of temperature and pressure conditions, as shown, for example at 623 K and 300 Torr in Fig. 5.
Figure 5: Representative high-temperature DCS spectrum, at 623 Kelvin, 300 Torr (brown crosses) [malarich_dual_2021]. Much of the HITRAN2020 residual (green dotted traces) is due to features that have the default $E"\textsubscript{HITRAN2020}=999.99$ cm-1. The net negative residual indicates that $E"\textsubscript{HITRAN2020}$ is too large. Orange trace shows final model with updates described here and in [malarich_dual_2021].
2.1.11 IAO & LiPhy CRDS measurements 7606–7919 cm-1
An extended assignment and analysis of 12CH4 lines in the 7606–7919 cm-1 region was recently reported by Nikitin et al. [nikitin_first_2024]. The experimental methane spectra in this range recorded with CRDS technique in Grenoble [mondelain_2011] at 80 K and 296 K correspond to the high-energy part of the WKLMC empirical line list [campargue_2013]. Some lines in this range belong to the upper edge of the Icosad band system previously considered in [rey_2015, rey_2016], but most of the other transitions have been included in HITRAN2020 without assignments.
Quantum assignments for 1382 transitions pointing to 33 vibrational sub-levels of the 12CH4 Triacontad system were obtained using a combined approach based on variational ab initio predictions [rey_2013, rey_new_2018] with a subsequent empirical optimization of the effective spectroscopic models derived from the potential and dipole moment energy surfaces as described in [tyuterev_2022, rey_2022]. They belong to three weak Triacontad bands, $6\nu_{4}$, $\nu_{2}$$5\nu_{4}$, $\nu_{1}$$4\nu_{4}$, falling in the methane transparency window near 1.28 µm. The positions of these vibrational bands are in good agreement with the values obtained from the ab initio potential energy surface [nikitin_var_2018, nikitin_pes_2025]. The root-mean-square fit deviation for the line positions of newly assigned transitions was 0.006 cm-1. The intensities of the line list for these bands span the range $10^{-25}$ to $9\times 10^{-30}$ cm/molecule. The full statistical results of the analyses are given in Table 1 of [nikitin_first_2024] and an overview plot of the 12CH4 experimental line lists near 1.28 µm is shown in Figure 4 of [nikitin_first_2024].
2.1.12 LiPhy Kitt peak FTS and CRDS measurements 10802–13922 cm-1
Figure 6: Update of the methane line list above 10800 cm-1. The HITRAN2020 list (orange circles) was completed by new FTS and CRDS measurements (blue and green circles, respectively). The TheoReTS variational list [rey_new_2018] (grey circles) is plotted as background below 13500 cm-1.
In Ref. [campargue_high_2023], empirical line lists at room temperature were retrieved from a long-path Kitt Peak FTS spectrum between 10802 and 13922 cm-1 (about 12800 lines) and from a high-sensitivity CRDS spectrum in the 13060–13300 cm-1 interval (about 2650 lines). Above 10800 cm-1, the previous HITRAN2020 dataset was limited to the 10923–11502 cm-1 interval where measurements obtained by intracavity laser absorption spectroscopy (ICLAS) were used as source (see Brown et al. [brown_methane_2013]). The FTS results of Ref. [campargue_high_2023] were found to be in good agreement with the HITRAN2020 dataset.
At these wavenumbers, the methane spectrum is especially highly congested even at room temperature, and, according to Ref. [campargue_high_2023], part of the absorption shows up as a “quasi-continuum”, which should be added to the absorption lines to fully account for the methane absorption in the region. The “quasi-continuum” represents about one third of the total absorption in the 11000–11400 cm-1 region. Note that the intensity sum of the lines and of the continuum was found to be in good agreement with the intensity predictions of the TheoReTS list [rey_new_2018].
2.2 Updates in HITRAN
The additions and updates from the data described in Section 2.1 is discussed here. As mentioned above, HITRAN2020 was not simply replaced by these lists provided to the HITRAN team. Instead, a series of validation processes involving comparisons with experimental spectra and previous parameters were used.
2.2.1 900–1050 cm-1
The new line list from Nikitin et al. [nikitin_improved_2024] (12CH4, 13CH4, 12CH3D & 13CH3D) completes the lower energy part of the dyad region, adding many lines including those for the lower-abundance 13CH3D isotopologue. Their data were merged directly to the region, replacing the transition wavenumbers for any difference ($\Delta\tilde{\nu}=0$ cm-1) between HITRAN2020 and the new parameters, the lower energy levels for $\Delta E^{\prime\prime}=0.1$ cm-1, and the intensities for $\Delta S_{ij}/S_{ij}=0.1$. About half of the lines in the list served mainly to update the positions and intensities with a few energy levels and assignments, while the other half was entirely new.
2.2.2 2500–3900 cm-1
The pentad region is relatively well defined in HITRAN2020, nevertheless, we have incorporated a few noticeable improvements. Spectra from Hargreaves et al. [2015ApJ...813...12H] ($S1$) and Bertin & Vander Auwera [bertin_co2_2024] ($S2$) were used to validate the spectral parameters. Although $S1$ spans a much larger region (2500–4900 cm-1), $S2$ (2860–3160 cm-1) provides a higher signal-to-noise ratio (SNR) and higher resolution, making it a better choice for validating the line list in this window. Small missing parts appear as well in $S1$ from cutting water absorption; making these small windows unusable for validation.
Figure 7: Residuals obtained at 296 K between experiment from Bertin & Vander Auwersa [bertin_co2_2024] and Voigt calculations from the line lists used for the HITRAN2024 update
The residuals between $S1$, $S2$ and the models calculated using the HITRAN Application Programming Interface (HAPI) with lines from ExoMol (12CH4), MeCaSDa (12CH4 & 13CH4), and HITEMP (12CH4 & 13CH4) were used to compare the quality of each parameter on a line-by-line basis (these residuals are presended for $S2$ in Fig. 7). Despite the good quality of fit, HITEMP was used when all other data were insufficient, since many lines are not fully assigned. By artificially increasing the pressure of the models, we could isolate the missing or strongly shifted lines in every line list, resulting in downward-pointing peaks. Using a combination of derivatives and thresholding, we could automatically find which areas were missing in HITRAN and which line lists could cover them. Because the residuals can be caused by multiple missing lines, every line above a certain intensity in a small window around the region to update was taken from the new line list. We then used the same technique, without the pressure increase, to find and update the weakly shifted lines in HITRAN. In addition to this automated approach, manual selections were made between lists for lines that were incorrectly updated or missed because of the various thresholds.
2.2.3 3900–4900 cm-1
Spectra $S1$ and from Birk et al. [birk_esa_2017] ($S3$), including CRDS spectra not shown here, were used for the octad region. For the same reason as above, $S3$ was used for its window of 3900–4500 cm-1 instead of $S1$.
Figure 8: Residuals obtained at 296 K and 0.17143 hPa between the experiment from Ref. [birk_esa_2017] and Voigt calculations from the line lists used for the HITRAN2024 update.
Figure 9: Residuals obtained at 198 K and 0.16941 hPa between the experiment from Ref. [birk_esa_2017] and Voigt calculations from the line lists used for the HITRAN2024 update.
Figure 10: Residuals obtained between experiment from Birk and Voigt calculations from the past and new HITRAN versions
Unlike in the pentad region, several new line lists presented a clear improvement over large portions of the spectrum, as illustrated in Figures 8 and 9. This result alleviated the need for line-by-line analysis, with a few exceptions that were then fixed. In those regions, the new lists completely replaced HITRAN2020, adding new lines and updating existing ones unless no significant changes were made by the conditions detailed in Sec. 2.2.1. The residuals were then analyzed, keeping HITRAN2020 parameters if needed, ensuring that only the best parameters remained. This method was used for the data from Rodina et al. [rodina_improved_2021], between 4100 and 4190 cm-1, improving both CH4, 13CH4, and CH3D. Their line list extended to 4300 cm-1, but that region was previously updated in HITRAN2020 for TROPOMI [hu_operational_2016]. A list provided by Sung et al. for 13CH4 has been added as well between 4000–4700 cm-1, including the region between 4190–4300 cm-1 which replaces some lines, or intensities, incorrectly assigned to the main isotopologues. The list from Nikitin et al. [nikitin_preliminary_2013] improving CH3D between 4000–4554 cm-1 has been added, adding the lines missing in HITRAN2016 [gordon_hitran2016_2017] when the preliminary version of their data was first used. Finally, the empirical list from Toon (atm.161) has been added between 4341–4900 cm-1, improving CH4, 13CH4, and CH3D.
We observed an approximately systematic shift in the transition wavenumber of lines from Daumont et al. [daumont_new_2013]. Applying a constant shift of 0.001 cm-1 appeared to improve the residuals, particularly below 4200 cm-1. The effect is limited above 4375 cm-1 as the addition of ATM list already improved that region appreciably. The final residuals for the latest HITRAN versions are shown in Figure 10, confirming the great improvement above 4375 cm-1, after keeping the region unchanged in the last 3 versions.
2.2.4 4970–5300 cm-1
The new data from Starikova et al. [starikova_assignment_2024] partially fills the gap of 13CH4 in the lower part of the tetradecad region. More than 1600 new lines were incorporated in HITRAN2024 (see Fig. 2). No 13CH4 lines were present in this region in HITRAN2020.
2.2.5 5900–6530 cm-1
Figure 11: Residuals obtained at 294 K and 0.5 hPa between the experiment from Ref. [birk_measurement_2025] and Voigt calculations from the line lists used for the HITRAN2024 update.
The DLR/NIST line list [reed_multilaboratory_2025, yin_2025] based on Nikitin 2017 [nikitin_analysis_2017] replaced the GOSAT 2014 12CH4 transition wavenumbers, and intensities in their region, for more than 2500 transitions. In several manifolds, better residuals were achieved with fewer lines than previously included in HITRAN2020 (Fig. 11). As such, some lines were removed to accommodate for these changes as well. The new pressure-induced parameters were used to replace the previous values, as discussed below in Sec. 3.
The new list from Ben Fathallah et al. [benfathallah_ch3d_2024] uses their experimental 12CH3D data, and variational calculations if not available, for inclusion in HITRAN2024. More than 5800 new transitions supplement the lower wavenumber range of the previous data. These data comprise 4800 lines providing updates on the positions, intensities, energy levels, as well as new assignments.
2.2.6 6770–7570 cm-1
The new PLDL list [malarich_dual_2021] mostly replaces the 12CH4 transition wavenumbers and some lower energy levels of HITRAN2020 while keeping the old intensities. More than 22,000 transitions were updated using their data, and about 300 new lines were added. Similarly to the DLR/NIST/USTC measurements, the new pressure-induced parameters were used as described in Section 3.
2.2.7 7600–7920 cm-1
The line list provided by Nikitin et al. [nikitin_first_2024] completes the 12CH4 main band in the lower part of the triacontad. More than 400 new lines were added, and more than 50 were updated on their positions, intensities and lower state energy levels. Another major benefit of the update in that region comes with the addition of rotational assignments, which were previously very scarce and are now mostly complete for these bands. With $E_{\mathrm{Lower}}$, this allows for modeling in a wider range of temperatures and gives better estimation of line-shape parameters.
2.2.8 10800-14000 cm-1
This new list from Campargue et al. [campargue_high_2023] completes the update below 10923 cm-1 and above 11502 cm-1: regions that were previously empty (see Fig. 6), with an additional 5000 FTS and 1800 CRDS new transitions, and excluding the quasi-continuum, which contributes to the total absorption and will need to be accounted for in the future. These lines are new to HITRAN. As a consequence of being completely empirical, the new lines were attributed to the main isotopologue by default. This issue will have to be addressed in a future HITRAN update.
2.3 Overview
In total, about 100,000 lines have either been modified or added from the lists described above to the database, increasing the total number of lines from 446,000 in HITRAN2020 to 485,000 now.
Figure 12: Line list overviews: Intensities of each line included in HITRAN2020 (upper panel) and HITRAN2024 (lower panel) for every isotopologues of methane
The extent of these additions can be visualized in Fig. 12, where we see that 12CH4 lines have been expanded up to 14000 cm-1 and, with the other isotopologues, completed some of the gaps. Additionally, 300 HITRAN2020 lines were removed; however, updated unassigned lines, or those with an incorrect assignment format in HITRAN2020, are counted as removed and added. These constitute the majority of removed lines, because they could not be automatically matched in other line lists. Some of them now have complete or at least partial assignments.
Table 2: Source of the new line parameters in the HITRAN2024 update
Line list Isotopologue
Transition wavenumbers
Intensities Region covered (cm-1)
Nikitin et al., 2025 [nikitin_improved_2024] 12CH4 1880 1610 900 – 1050
13CH4 60 40
12CH3D 500 460
13CH3D 60 60
ExoMol [tennyson_2024_2024] 12CH4 550 450 2503 – 4678
MeCaSDa [RICHARD2024109127] 12CH4 70 30 2786 – 4532
13CH4 40 30
Sung et al. [sung_new_2024] 13CH4 5680 5650 4000 – 4700
Nikitin et al., 2013 [nikitin_preliminary_2013] 12CH3D 690 670 4001 – 4537
Rodina et al., 2021 [rodina_improved_2021] 12CH4 4930 4220 4100 – 4190
13CH4 430 430
12CH3D 10 10
atm.161 12CH4 22410 23350 4340 – 4899
13CH4 770 790
12CH3D 140 140
Starikova et al., 2024 [starikova_assignment_2024] 13CH4 1640 1640 4970 – 5300
Reed et al., 2025 [reed_multilaboratory_2025] 12CH4 3020 2710 6016 – 6115
Ben Fathallah et al., 2024 [benfathallah_ch3d_2024] 12CH3D 10640 5850 6098 – 6530
Malarich et al., 2021 [malarich_dual_2021] 12CH4 22320 280 6770 – 7570
Nikitin et al., 2024 [nikitin_first_2024] 12CH4 480 480 7629 – 7918
Campargue et al., 2023 [campargue_high_2023] 12CH4 6780 6780 10803 – 13923
In general, more transition wavenumbers were taken from the new data than line intensities (see Table 2), which stems from our criteria for replacing the existing parameters. From HITRAN2020, about 56,000 of the transition wavenumbers were replaced, with 16,000 for the intensities, 7500 for the energy levels, and 3400 for the assignments. Given the general lower accuracy of the intensities provided, a larger difference between the HITRAN2020 values and the new data was required before updating the value. In most cases, uncertainties were not even reported. The SNR is another contributing factor to consider when most residuals at room temperature and natural isotopic abundance do not always exhibit an obvious improvement from the new lists, especially for weaker lines.
3 Pressure-induced line-shape parameters
3.1 USTC, NIST & DLR measurements
Figure 13: An overview of CRDS spectra and global fit results for the R(0)-R(10) transition with LMVP.
[…]
For transitions in the $2\nu_{3}$ band of methane, HITRAN2020 only covers the parameters of Voigt profile (VP), i.e., pressure broadening ($\gamma_{0}$) and pressure shifting ($\delta_{0}$). Yet, to achieve greater precision in methane atmospheric simulations, more intricate line shapes and precise parameters are needed. In this regard, we employed the Speed-Dependent Voigt profile (SDVP) corrected for first-order Line-Mixing (LM), abbreviated as LMSDVP, as our base model to analyze the CRDS data. SDVP is a degregation of the pCqSDHC(HT) profile discussed in Ngo et al.[NGO2013JQSRT], which is more accurate than the VP owing to the inclusion of the speed-dependent effect. The involvement of numerous fitting parameters in the HT profile, compounded by the intricate nature of methane spectra, can result in closely numerically correlated spectral line parameters. This can impede the convergence of fitting solutions when analyzing the actual spectra or cause the derived physical parameters to lack physicality. Therefore, we did not use the HT profile. Indeed, the SDVP is adequate for most endeavors of spectral analysis and simulations, but additional treatment is needed due to the high degeneracy of methane states. Thus we introduced the first-order Rosenkranz parameter parameter, $Y_{n}$, to account for line mixing effects [Rosenkranz1975]. From a computational perspective, $Y_{n}$ exhibits a strong correlation with line intensity $S$. Consequently, an independent measurement of line intensity is crucial for acquiring dependable line-mixing parameters. Together with the intensity measurements by NIST and DLR, a global-fit procedure was developed based on the LMSDV model. We also converted the measured nitrogen-broadening coefficients to air-broadening coefficients using the formula $\gamma_{0}^{\text{air}}=0.79\gamma_{0}^{\mathrm{N}_{2}}+0.21\gamma_{0}^{\mathrm{O}_{2}}$, and utilizing polynomial parameters for $\gamma_{0}^{\mathrm{O}_{2}}$ as provided by Lyulin et al. [Lyulin2009JQSRT]. The nitrogen pressure shifting coefficients were similarly scaled as $\delta_{0}^{\text{air}}=0.79\delta_{0}^{\mathrm{N}_{2}}+0.21\delta_{0}^{\mathrm{O}_{2}}$, again utilizing parameters for $\delta_{0}^{\mathrm{O}_{2}}$ as provided by Lyulin et al. [Lyulin2009JQSRT]. Values of $Y_{n}^{\mathrm{N\textsubscript{2}}}$ for the measured CH4–N2 system are converted for CH4–air through multiplication by 0.985, as shown in Tran et al. for the $2\nu_{3}$ band [Tran2010].
Many weaker transitions from 5970 – 6250 cm-1, outside the NIST/DLR/USTC $2\nu_{3}$ overlap set, were updated with Voigt lineshape parameters ($\gamma_{0},\delta_{0}$) recently determined from ambient temperature (nominally 298 K) FTS measurements performed by Birk et al. at DLR. The initial line list was the Nikitin 2017 line list [nikitin_analysis_2017], which was based on the empirical WKLMC line list, updated with the line intensity parameters described in Reed et al. [reed_multilaboratory_2025] and the line positions of Votava et al. [votava_2022], where available. Improved Voigt parameters are reported for approximately 1650 transitions with $S>3\times 10^{-23}$ cm-1/(molec.cm-2) outside the $2\nu_{3}$ band. Lineshape parameters for $2\nu_{3}$ transitions were floated in the fit; however, the values adopted in HITRAN2024 are those determined by USTC above.
3.2 12CH4 & 13CH4
Broadening parameters ($\gamma_{0}$) were gathered from all post-year 2000 sources [antony_n2_2008, bertin_co2_2024, clark_difference_2004, devi_self-_2015, devi_spectral_2016, devi_multispectrum_2018, es-sebbar_intensities_2014, es-sebbar_linestrengths_2021, farji_airinduced_2021, gabard_line_2010, gharib-nezhad_h2induced_2019, ghysels_spectroscopy_2011, ghysels_temperature_2014, grigoriev_estimation_2001, lyulin_measurements_2014, ma_temperature_2016, manne_determination_2017, mondelain_line_2005, mondelain_measurement_2007, pine_speeddependent_2019, pine_speeddependent_2000, pine_multispectrum_2003, richard_self_2023] that we could find, and separated by collision partners, transition symmetries, and ranking index. Unlike for other rotors, only the ranking index, $\alpha$, lets us differentiate between lines of the same manifold and symmetry, which simply relates to the energy level instead of a motion. Because of this, the parameters do not exhibit any clear pattern with $m$ for each index (see Fig. 14). Thus, separation was only used to isolate the most intense lines, which are found in $\alpha=1$.
Figure 14: Air-broadening of methane for the first ranking index, $\alpha^{\prime\prime}=1$, displayed from the many sources available in literature. The HITRAN2020 data (in lighter blue) is shown with transparency. The more opaque points show multiple transitions having the same value.
The different line profiles used in these studies impact the value of the retrieved coefficients by a small but non-negligible amount. For consistency with the parameters already present in HITRAN, we chose to focus on the Voigt profile. Separating each profile would greatly reduce the amount of data per fit; therefore, we scaled the respective non-Voigt parameters by the average Voigt to non-Voigt ratio for each transition. This approach could only be performed in cases where data were fit with both non-Voigt and Voigt profiles —thus not all studies could be included. To further increase the amount of available data, the CH4–N2 pressure broadening parameters were combined with those of –air, and scaled by a factor of 0.988 taken from Pine & Gabard [pine_speeddependent_2000]. This approach was previously used in other studies [mondelain_measurement_2007] and agrees very well with other measurements (see Fig. 15). Finally, thanks to the weak branch dependence, broadening parameters for the $P$, $Q$, and $R$ branches were combined. Ideally each branch should be separate, however, the broadening measurements in methane are too uncertain for the difference to be reliably quantified.
Using a few assumptions, we can constrain the Padé approximant function to ensure convergence of the fit. First, because pressure broadening is positive-valued, we will force every parameter to be positive for simplicity. This constraint also prevents discontinuities in the curve by removing the possibility of division by zero. Second, based on measurements in various species, not only methane, we can assume that in the $\lim_{J\to\infty}$, $\gamma_{0}$ is equal to zero. For this reason, it is preferable to have a polynomial of greater order in the divisor. We found that a [3/4] function of the form
$\gamma_{0}=\dfrac{a_{0}+a_{1}m+a_{2}m^{2}+a_{3}m^{3}}{1+b_{1}m+b_{2}m^{2}+b_{3}m^{3}+b_{4}m^{4}}$ (1)
works well for our level of precision. As observed in Fig. 14, even a single symmetry and ranking index does not lead to a smooth looking curve. Some transitions give systematically stronger or weaker parameters. Good examples of that behavior are the $J^{\prime\prime}=4$, $\alpha=1$ transitions of the $F_{2}$ rotational symmetry, both in $P$ and $R$ branches, with parameters being weaker than those of their surrounding lines. This behavior appears in most experiments and is probably not an experimental artifact. Because these values deteriorate the fit, such points are removed beforehand.
Figure 15: Data used for the fit, and resulting Padé approximant in the $F_{2}$ symmetry, combining $P$ and $R$ branches
Figure 16: Air-broadening data used for the fit, and resulting Padé approximant for all symmetries
The results (Fig. 16) yield only a single function per symmetry. However, given the lack of data at higher quanta and ranking indices, these are the only reliable Padé curves that could be calculated. Nevertheless, they should still provide a good consistent set of parameters for spectroscopic observations. The $A_{1}$ symmetry diverges rapidly from the others. This could simply be a problem with the fit, but again, no assumption at higher quanta should be made, therefore, we decided to keep it. The choice of parameters added to HITRAN2024 was based on the data available. Not every transition used the Padé value. Specifically, if high-precision measurements were available, they were taken instead. This rule holds not only for new data, but those present in HITRAN2020 as well, which were not replaced if the uncertainty is lower. With this approach, we are able to conserve the structures observed in the experiments. The air-broadening half-widths of the $\nu_{3}$ band from HITRAN2020 were replaced using this rule, ensuring a traceable set of data for future updates. Some HITRAN values were found to be unrealistically low or high, possibly caused by typographical errors, and were thus replaced as well. Using this approach, the total number of replaced broadening parameters was 30,000.
3.3 12CH3D & 13CH3D
With its different symmetry, CH3D must be treated differently, and benefits from a more rigorous separation based on quantum numbers. A better semi-empirical approach developed by Dudaryonok et al. [dudaryonok_semiempirical_2018] seems to produce very good results in the $\nu_{6}$ band. Ignoring the wavenumber dependence, we applied this method for the broadening parameter of every new line added in Sec. 2. The wavenumber dependence of the pressure shift was reintroduced, summing the semi-empirical value to the empirical formula in HITRAN2012 [brown_methane_2013] of $2\times 10^{-6}\times\Delta\nu$, limited to negative values. A similar approach was used where the extremely low or high values of HITRAN2020 were replaced along with the new lines, which only amount to about 70.
4 Conclusion
The parameters of more than 58,000 lines of 12CH4, 13CH4, and 12CH3D have been updated for HITRAN20204 using the new line lists provided. Transitions wavenumbers, intensities, and lower-state energy levels were taken directly from these lists to replace the values from HITRAN2020. Laboratory spectra in the pentad and octad regions were used by the HITRAN team to validate and choose between the overlapping data sets, while other regions used the validations from the other groups. 40,000 lines of the various isotopologues have been added, including some lower-intensity lines that were previously missing. In regions that were previously empirically determined, this work allowed us to reassign some of the intensities to the correct isotopologues. Assignments of many lines have been completed between the octad and triacontad regions.
The broadening parameters of these newly added lines were replaced using Padé-approximant functions fit from multiple recent studies for 12CH4 and 13CH4, unless reliable experimental data were provided, such as those generated at the USTC. The functions were obtained from data sets separated based on their rotational symmetry and ranking index. Currently, only one function per symmetry was used, regardless of the ranking index. The pressure-induced shifts were calculated using the same formula from HITRAN2012 [brown_methane_2013]. Semi-empirical functions from Dudaryonok et al. [dudaryonok_semiempirical_2018] were used for 12CH3D. The pressure shifts were scaled using the HITRAN2012 formula to account for the vibrational dependence, while the broadening parameters were taken as is. The broadening parameters from the $\nu_{3}$ band, were replaced similarly using the same Padé approximants, however, experimental values were used when available.
[…]
The relational structure of the HITRAN database [Hill2016] and functionality of HAPI [Kochanov2016] allow one to add non-Voigt line shape parameters. For instance, in this edition, we have added the speed-dependent Voigt parameters and first-order line mixing for air-broadening in the tetradecad region. It will be important to have a consistent set of parameters in all the bands as was done for CO2 [Hashemi2020], N2O, and CO [Hashemi2021] in HITRAN2020. Of course, it is much harder for methane as there are strong symmetry and vibrational dependencies in comparison with linear molecules. Moreover, it will be important to introduce the full line mixing parameters, where possible. The accuracy of first-order line mixing deteriorates quickly at pressures approaching $1$ atm and above. Dense spectral regions such as $Q$ branches are especially affected and need full line mixing, possibly in combination with speed-dependent effects [ciurylo_speeddependent_2000, bertin_co2_2024], to be properly measured. Unlike some other parameters, these require complete knowledge of the assignments to be implemented; therefore, more theoretical work will have to be done in preparation. Currently, data from Tran et al. is available for CH4–air [Tran2006, Tran2010] and other collision partners [tran_model_2006-1, Tran2022], and will be considered in future updates.
In addition, in order to support the studies of planetary and brown dwarf atmospheres, it is important to provide parameters associated with broadening by “planetary” gases, including H2O [Tan2019], H2, He and CO2 [Tan2022]. Unlike many other gases for which these “planetary” broadening parameters are available, only broadening by water vapor is available for methane. In this work, we have used the scaling factors recommended in Ref. [Tan2019] to assign broadening by water vapor values to all new lines of methane. We have collected existing information on broadening by hydrogen, helium, and carbon dioxide (including Refs. [bertin_co2_2024, Tran2022, Varanasi1971, Varanasi1974, Varanasi1990, Fox1989, Gabard2013, es-sebbar_intensities_2014, gharib-nezhad_h2induced_2019, Vispoel2019, Sung2020, Yousefi2021, Hosokawa2025, Clment2025], but it is still a lot of work to process these data, which often do not agree with each other within stated uncertainties.
Author contributions
T. Bertin: Methodology, Software, Validation, Investigation, Resources, Data Curation, Writing – Original Draft, Writing – Review & Editing, Visualization. I. E. Gordon: Conceptualization, Methodology, Investigation, Resources, Data Curation, Writing – Original Draft, Writing – Review & Editing, Supervision, Project administration, Funding acquisition. R. J. Hargreaves: Resources, Data Curation, Writing – Original Draft, Writing – Review & Editing. J. Tennyson: Resources, Writing – Original Draft, Writing – Review & Editing. S. N. Yurchenko: Resources, Writing – Original Draft, Writing – Review & Editing. K. Kefala: Resources. V. Boudon: Resources, Writing – Original Draft. C. Richard: Resources, Writing – Original Draft. A. V. Nikitin: Resources, Writing – Original Draft. V. G. Tyuterev: Resources, Writing – Original Draft, Writing – Review & Editing. M. Rey: Resources, Writing – Original Draft, Writing – Review & Editing. M. Birk: Resources. G. Wagner: Resources. K. Sung: Resources, Writing – Original Draft, Visualization. B. P. Coy: Resources. W. Broussard: Resources. G. C. Toon: Validation, Investigation, Resources, Writing – Review & Editing. A. A. Rodina: Resources. E. Starikova: Resources, Writing – Original Draft, Visualization. A. Campargue: Resources, Writing – Original Draft, Writing – Review & Editing, Visualization. Z. D. Reed: Resources, Writing – Original Draft, Writing – Review & Editing, Visualization. J. T. Hodges: Resources, Writing – Original Draft, Writing – Review & Editing. Y. Tan: Resources, Writing – Original Draft, Writing – Review & Editing, Visualization. N. A. Malarich: Resources. G. B. Rieker: Resources, Writing – Original Draft, Visualization.
Acknowledgments
The update of the methane line list in HITRAN is funded through NASA grants 80NSSC23K1596, and 80NSSC24K0080. Work at UCL was supported by ERC Advanced Investigator Project 883830; we thank Jingxin Zhang for help generating the HITRAN-style MM line list.
The research at the Jet Propulsion Laboratory, California Institute of Technology, was performed under contract with the National Aeronautics and Space Administration. K. Sung acknowledges JPL Summer Internship Program support for B. P. Coy and A. W. Broussard. Partial support from NASA-ROSES-XRP and from RSF-ANR TEMMEX Russian-French project. Research at the National Institute of Standards and Technology was supported in part by the NIST Greenhouse Gas and Climate Science Program.
Disclaimer
Identification of certain commercial equipment, instruments, software, or materials does not imply recommendation or endorsement by the National Institute of Standards and Technology, nor does it imply that the products identified are necessarily the best available for the purpose.
Appendix A Abbreviations
The following abbreviations and acronyms have been used throughout the article:
  • 1.
    CRDS – Cavity ring-down spectroscopy
  • 2.
    DLR – Deutsches Zentrum für Luft und Raumfahrt (German Aerospace Center)
  • 3.
    FTIR – Fourier transform infrared
  • 4.
    FTS – Fourier transform spectrometer
  • 5.
    GOSAT – Greenhouse Gases Observing Satellite
  • 6.
    HAPI – HITRAN Application Programming Interface
  • 7.
    HITRAN – High-resolution transmission molecular absorption database
  • 8.
    HITEMP – High-temperature molecular spectroscopic database
  • 9.
    HT – Hartmann–Tran
  • 10.
    IAO – Institute of Atmospheric Optics
  • 11.
    ICB – Laboratoire de l’Université de Bourgogne
  • 12.
    JPL – Jet Propulsion Laboratory
  • 13.
    LM – Line mixing
  • 14.
    LMSDVP – Line-mixing speed-dependent Voigt profile
  • 15.
    MARVEL – Measured Active Rotational-Vibrational Energy Levels
  • 16.
    MeCaSDa – Methane Calculated Spectroscopic Database
  • 17.
    NASA – National Aeronautics and Space Administration
  • 18.
    NIST – National Institute of Standards and Technology
  • 19.
    pCqSDHC – partially-Correlated quadratic-Speed-Dependent Hard-Collision
  • 20.
    PES – Potential energy surface
  • 21.
    SDVP – Speed-dependent Voigt profile
  • 22.
    SEOM-IAS – Scientific Exploitation of Operational Missions — Improved Atmospheric Spectroscopy Databases
  • 23.
    SNR – Signal to noise ratio
  • 24.
    TCCON – Total Carbon Column Observing Network
  • 25.
    TheoReTS – Theoretical Reims-Tomsk Spectral data
  • 26.
    TROPOMI – Tropospheric Monitoring Instrument
  • 27.
    UCL – University College London
  • 28.
    VAMDC – Virtual Atomic and Molecular Data Centre
  • 29.
    VP – Voigt profile